# 1]FAIR at Meta Dualformer: Controllable Fast and Slow Thinking by Learning with Randomized Reasoning Traces

### v2

@@ 5.1.1 Fast Mode
Dualformer also consistently generates more diverse plans for all the considered problems. In Appendix 9, we show one Maze example and plot the unique correct plans generated by Dualformer and the baseline model. One interesting observation is that the diversity score of Dualformer, i.e., the average number of unique correct plans out of 64 responses, increases as the Maze size goes up. Intuitively, as the Maze becomes larger, there are more possible routes to reach a single goal location.
This suggests that Dualformer learns the Maze structure, while the Solution-Only model is potentially memorizing the optimal plans as its diversity score is close to 1 for all the Maze sizes.
@@ 5.1.2 Slow Mode
Table 5.3 reports the results when Dualformer operates in slow mode. The corresponding baseline is the Complete-Trace model, which uses the same architecture and is trained on data with complete A* search traces. In addition to the metrics reported before, we report the average length of reasoning traces across the 64 responses, aggregated over all the 1000 evaluation tasks. The results show that Dualformer achieves both enhanced planning power and reasoning speed. It outperforms the Complete-Trace model for all the correctness and optimality metrics: solved rates, optimal rates, and SWC. Moreover, the reasoning trace yielded by Dualformer is notably shorter than the baseline model. On average, Dualformer reduces the trace length by 49.4% across the five tasks. As before, Dualformer also generates more diverse plans compared with the baseline. We refer the readers to Appendix 9 for concrete examples.
@@ 5.1.5 Ablation Study
Figure 10.1 compares the 1-Optimal-64 rate of Dualformer and Mix- models. We test 8 values of : (equivalent to the Complete-Trace model), , and (equivalent to the Solution-Only model). In both inference modes, Dualformer beats all the Mix- models for all the five tasks we consider. In fact, Dualformer also outperforms Mix- models in the other metrics we consider too, see Section 10. In particular, it is the “fastest” one when operating in slow mode: it generates the shortest reasoning trace.
@@
We compare Dualformer with its variants where we halt the dropping strategies at specific levels. For the Maze environments, we fix the probability of using the complete reasoning trace for a training example to be for all the variants and vary the other probabilities. For the Sokoban environments, we fix the probability of Level 1 dropping to be . Table 11.1 lists the probabilities of different dropping strategies we use for all the models, and Table 5.7 shows the results. It should be noted that the variants cannot operate in fast mode, since they are not trained on any Level 4 data. Therefore, we only report the slow mode performance. As we increase the strategy level, the length of the reasoning trace decreases. In regard to the other metrics, the performance of Level 1+2+3 model and Dualformer are comparable. However, Dualformer enjoys the advantage of shorter reasoning trace and the capability to function in fast mode.
@@ 5.2 Application to LLM Training: Math Reasoning
We evaluate all the models using a dataset named Aug-MATH, which is derived from the MATH dataset Hendrycks et al. (2021) that contains 7500 training examples of math questions and solutions, and 5000 testing examples. Following Yu et al. (2023), we query the Llama-3.1-70B-Instruct model to rewrite the solutions to include more detailed intermediate steps, following a given format. To encourage the diversity of reasoning traces, we sample 4 LLM responses for each problem using a temperature of 0.7 and top-. The resulting dataset then contains training examples and testing examples. In Appendix 12.1, we show the prompt template that we use for solution rewriting, and a concrete training example before and after rewriting.
@@
We finetune two base models Mistral-7B and LLama-3-8B for two epochs using a batch size of . We use the AdamW optimizer Loshchilov and Hutter (2019) with a learning rate of for the Mistral models, and for the Llama-3 models.
@@
Similar to Dualformer, we evaluate the models in both the fast and the slow mode, where the LLM is required to output the final solution directly or to solve the problem step by step. Following Yu et al. (2023), we use zero-shot prompting when evaluating the models. See Appendix 12.2 for the prompts we use in each mode. We consider the Greedy@1 metric Dubey et al. (2024); Yu et al. (2023), where for each question we generate 1 response using a temperature of 0 and verify the correctness. We also report the pass@20 metric Chen et al. (2021), where we randomly sample 20 responses using a temperature of 0.5.
For reference, we also report the results of fine-tuning those models on the original MATH dataset.
@@ Results
The results are presented in Table 5.8. We test four values of : and . Our results show that the proposed training strategy makes both LLMs more effective and efficient. We first inspect the results of the Mistral-7B models. For the slow mode inference, fine-tuning the model with trace dropping and randomized training improves upon the baseline model that is directly finetuned on Aug-MATH dataset. The absolute Greedy@1 metric improved by 1.7% when (which amounts to 10% relative performance improvement), and 0.9% with and , and 0.1% when .
Our models are also outperforming the baseline model for the Pass@20 metric when and , where the absolute correct rate increases to . Under both evaluation schemes, the average length of the reasoning trace goes down as goes up.
Similarly, for inference in the fast mode, our models also achieve higher correct rate. A similar trend of performance improvement also holds for the Llama-3-8B models. Finally, for reader’s reference, we include the results of both the Mistral-7B and the Llama-3-8B models finetuned on the original MATH dataset. The performance of those models clearly lags behind222The trace generated by models trained on MATH is much shorter, because the answers in MATH do not contain as many intermediate steps as those in Aug-MATH..
@@ 6 Conclusion
We present a simple and easy-to-implement framework for training Transformers to solve reasoning and planning tasks. We carefully probe the structure of the reasoning traces and design corresponding dropping strategies that imitate the shortcuts in human thinking process. By randomly applying the dropping strategies to training examples, the resulting model, Dualformer, can be controlled to execute in either fast or slow reasoning mode, or in auto mode where it decides the mode to engage. Dualformer achieves enhanced performance for the maze navigation tasks and the Sokoban game, while reducing the number of reasoning steps required. Remarkably, our approach is not limited to training task-specific models from scratch. We apply techniques in the same spirit to finetune LLMs to answer math questions and obtain improved performance. Last, the proposed framework also reduces computation consumption as the input sequences are shortened after trace dropping.
Future work could investigate whether our approach helps models scale, and explore methodologies such as curriculum learning and hierarchical planning to adapt Dualformer for more complex tasks.

### v3

@@ 4.1.1 Fast Mode
Dualformer also consistently generates more diverse plans for all the considered problems. In Appendix E, we show one Maze example and plot the unique correct plans generated by Dualformer and the baseline model. One interesting observation is that the diversity score of Dualformer, i.e., the average number of unique correct plans out of 64 responses, increases as the Maze size goes up. Intuitively, as the Maze becomes larger, there are more possible routes to reach a single goal location. This suggests that Dualformer learns the Maze structure, while the Solution-Only model is potentially memorizing the optimal plans as its diversity score is close to 1 for all the Maze sizes.
@@ 4.1.2 Slow Mode
Table 4.3 reports the results when Dualformer operates in slow mode. The corresponding baseline is the Complete-Trace model, which uses the same architecture and is trained on data with complete A* search traces. In addition to the metrics reported before, we report the average length of reasoning traces across the 64 responses, aggregated over all the 1000 evaluation tasks. The results show that Dualformer achieves both enhanced planning power and reasoning speed. It outperforms the Complete-Trace model for all the correctness and optimality metrics: solved rates, optimal rates, and SWC. Moreover, the reasoning trace yielded by Dualformer is notably shorter than the baseline model. On average, Dualformer reduces the trace length by 49.4% across the five tasks. As before, Dualformer also generates more diverse plans compared with the baseline. We refer the readers to Appendix E for concrete examples.
@@ 4.1.5 Ablation Study
Figure F.1 compares the 1-Optimal-64 rate of Dualformer and Mix- models. We test 8 values of : (equivalent to the Complete-Trace model), , and (equivalent to the Solution-Only model). In both inference modes, Dualformer beats all the Mix- models for all the five tasks we consider. In fact, Dualformer also outperforms Mix- models in the other metrics we consider too, see Appendix F. In particular, it is the “fastest” one when operating in slow mode: it generates the shortest reasoning trace.
@@
We compare Dualformer with its variants where we halt the dropping strategies at specific levels. For the Maze environments, we fix the probability of using the complete reasoning trace for a training example to be for all the variants and vary the other probabilities. For the Sokoban environments, we fix the probability of Level 1 dropping to be . Table G.1 lists the probabilities of different dropping strategies we use for all the models, and Table 4.7 shows the results. It should be noted that the variants cannot operate in fast mode, since they are not trained on any Level 4 data. Therefore, we only report the slow mode performance. As we increase the strategy level, the length of the reasoning trace decreases. In regard to the other metrics, the performance of Level 1+2+3 model and Dualformer are comparable. However, Dualformer enjoys the advantage of shorter reasoning trace and the capability to function in fast mode.
@@ 4.2 Application to LLM Training: Math Reasoning
We evaluate all the models using a dataset named Aug-MATH, which is derived from the MATH dataset Hendrycks et al. (2021) that contains 7500 training examples of math questions and solutions, and 5000 testing examples. Following Yu et al. (2023), we query the Llama-3.1-70B-Instruct model to rewrite the solutions to include more detailed intermediate steps, following a given format. To encourage the diversity of reasoning traces, we sample 4 LLM responses for each problem using a temperature of 0.7 and top-. The resulting dataset then contains training examples and testing examples. In Appendix H.1, we show the prompt template that we use for solution rewriting, and a concrete training example before and after rewriting.
@@
We finetune two base models Mistral-7B and LLama-3-8B for two epochs using a batch size of . We use the AdamW optimizer Loshchilov & Hutter (2019) with a learning rate of for the Mistral models, and for the Llama-3 models. Full training details, including how we select the learning rates and other hyperparameters, are deferred to Appendix C.2. For fine-tuning the models, we use the same prompt as in Yu et al. (2023), which is displayed in Appendix H.2.
@@
Similar to Dualformer, we evaluate the models in both the fast and the slow mode, where the LLM is required to output the final solution directly or to solve the problem step by step. Following Yu et al. (2023), we use zero-shot prompting when evaluating the models. See Appendix H.2 for the prompts we use in each mode. We consider the Greedy@1 metric Dubey et al. (2024); Yu et al. (2023), where for each question we generate 1 response using a temperature of 0 and verify the correctness. We also report the pass@20 metric Chen et al. (2021), where we randomly sample 20 responses using a temperature of 0.5. For reference, we also report the results of fine-tuning those models on the original MATH dataset.
@@ Results
The results are presented in Table 4.8. We test four values of : and . Our results show that the proposed training strategy makes both LLMs more effective and efficient. We first inspect the results of the Mistral-7B models. For the slow mode inference, fine-tuning the model with trace dropping and randomized training improves upon the baseline model that is directly finetuned on Aug-MATH dataset. The absolute Greedy@1 metric improved by 1.7% when (which amounts to 10% relative performance improvement), and 0.9% with and , and 0.1% when . Our models are also outperforming the baseline model for the Pass@20 metric when and , where the absolute correct rate increases to . Under both evaluation schemes, the average length of the reasoning trace goes down as goes up. Similarly, for inference in the fast mode, our models also achieve higher correct rate. A similar trend of performance improvement also holds for the Llama-3-8B models. Finally, for reader’s reference, we include the results of both the Mistral-7B and the Llama-3-8B models fine-tuned on the original MATH dataset, which clearly lags behind222The trace generated by models trained on MATH is much shorter, because the answers in MATH do not contain as many intermediate steps as those in Aug-MATH..
@@ 5 Conclusion
We present a simple and easy-to-implement framework for training Transformers to solve reasoning and planning tasks. We carefully probe the structure of the reasoning traces and design corresponding dropping strategies that imitate the shortcuts in human thinking process. By randomly applying the dropping strategies to training examples, the resulting model, Dualformer, can be controlled to execute in either fast or slow reasoning mode, or in auto mode where it decides the mode to engage. Dualformer achieves enhanced performance for the maze navigation tasks and the Sokoban game, while reducing the number of reasoning steps required. Remarkably, our approach is not limited to training task-specific models from scratch. We apply techniques in the same spirit to fine-tune LLMs to answer math questions and obtain improved performance. Last, the proposed framework also reduces computation consumption as the input sequences are shortened after trace dropping. Future work could investigate whether our approach helps models scale, and explore methodologies such as curriculum learning and hierarchical planning to adapt Dualformer for more complex tasks.
Appendix A Preliminaries
Our work builds upon the work of Lehnert et al. (2024). To perform planning, we train a Transformer to model a sequence of tokens that sequentially represents the planning task, the computation of A* algorithm, and the optimal solution derived from the A* search. The tokenization method is illustrated in Figure A.1. As a toy example, we consider a navigation task in a Maze where the goal is to find one shortest path from the start cell to goal cell, without hitting a wall cell. The A* algorithm has successfully determined an optimal plan. We use a sequence of tokens to express both the task and the Maze structure, which also serves as the prompt for Dualformer. The solution is depicted by the plan token sequence that describes the path using coordinates. The A* algorithm generates a search trace sequence that records the search dynamics performed, displayed in Figure 3.1. Recall that the A* algorithm is a pathfinding algorithm on a weighted graph. The create clause adds the node (represented by the subsequent coordinates) into the search frontier, and the close clause adds the node to the closed set. Each clause, either create or close, is followed by the tokens x, y, c0, and c1, representing the node’s coordinates, cost-since-start value, and heuristic values, respectively. For details of A* and the tokenization approach, we refer the readers to Russell & Norvig (2021) and Lehnert et al. (2024), respectively.
Appendix B Comparison of Fast Thinking and Slow Thinking
Appendix C Network Architecture and Hyperparameters
We use the same encoder-decoder Transformer architecture as in Lehnert et al. (2024) for Dualformer. It first converts each token into a one-hot vector, which is then transformed into a set of vectors through the embedding layer. The embedding vectors then go through the subsequent layers shown in Lehnert et al. (2024, Figure 4). We use RoPE embeddings for positional encoding, and no dropout is used in our architecture.
For the model size, architecture parameters, batch size, we use the same setup as in as Lehnert et al. (2024). Specifically, for the Maze tasks, we use a 15M-parameter model which consists of 3 attention heads and 6 layers with hidden size 64. We optimize the model by the AdamW (Loshchilov & Hutter, 2019) optimizer with learning rate 2.5e-4 and batch size 16, where and set to 0.9 and 0.99, respectively. A linear warm-up strategy was employed for the first 2000 gradient updates. Afterward, we use the cosine learning rate scheduler (Loshchilov & Hutter, 2016).
For the Sokoban tasks, we use a 46M-parameter model which consists of 4 attention heads and 8 layers with hidden size 96. We use batch size 64 and learning rate 7.5e-5, while the other hyperparameters are the same as above.
C.1 Environment & Dataset
We used the same dataset as in Lehnert et al. (2024), which is available at https://github.com/facebookresearch/searchformer. The Maze dataset contains examples and the Sokoban dataset contains examples and we use the first example sorted in accordance to their ”id” field in mongodb (following same approach as Lehnert et al. (2024)). For reader’s reference, the dataset is generated as follows. For the Maze tasks, 30-50% of the cells were first randomly designated as walls. Next, a start and a goal location were chosen randomly. Then, the algorithm was applied to generate an optimal plan. For the Sokoban tasks, we use a grid map where two wall cells were randomly inserted as obstacles. Moreover, two docks, two boxes, and two worker locations were placed randomly. Once a game is generated, it is only added to the dataset if it could be solved by the algorithm.
C.2 Math Reasoning
We use the implementation provided at https://github.com/meta-llama/llama-recipes for fine-tuning the models.
We train all the models for 2 epochs, using a batch size of 32. We use the AdamW optimizer with a learning rate of for the Llama model and for the Mistral models. The learning rate is selected as follows. We sweep over 3 values and choose the learning rate that yields the lowest validation loss. We then retrain the models using the selected learning rates on the full training dataset and report the results. We use the default values for the other hyperparameters. More specifically, we do not use linear rate warmup, weight decay, nor multistep gradient accumulation. We use betas=, eps=, (multiplicative step-wise learning rate decay) for AdamW and “packing” as the batching strategy.
Appendix D Controllable Generation of Dualformer
Dualformer offers flexible output options for users to choose. In this section, we demonstrate this by an example navigation task in a maze, see Figure D.1. The start location is (9, 10) and the goal location is (3, 6). The location coordinates and the maze structure is encoded in the original prompt below.
To configure Dualformer’s operation mode, we only need to append bos and a control token to the original prompt. To output solution only (fast mode), we inject plan. On the other hand, we inject create to let Dualformer output both trace and solution.
D.1 Fast Mode
The modified prompt for fast mode is displayed below. The extra tokens are highlighted in purple.
The following box shows the generated tokens, and Figure 1(a) plot the corresponding path.
D.2 Slow Mode
Similarly, we present the modified prompt for slow mode below. Comparing with the fast mode prompt, the only change is the control token becomes create.
The following box shows the generated output, and Figure 1(b) plots the corresponding trace and final path.
Appendix E Diversity of Generated Plans
The Dualformer outperforms baseline models in discovering unique feasible solutions. To illustrate this visually, we select one example maze task and generate 64 responses using Dualformer in fast mode. Figure E.1 plots all the unique feasible paths discovered by fast mode Dualformer alongside those found by the Solution-Only baseline (64 responses). Dualformer (fast mode) identified 42 unique feasible paths, while the Solution-Only model only found 3. Similarly, Figure E.2 compares slow mode Dualformer and the Complete-Trace (Searchformer) baseline. Dualformer (slow mode) discovered 39 unique feasible paths, whereas the Complete-Trace model only found 17.
Appendix F Comparison with Mix- Models
| Method | Avg Trace Length | 1-Optimal-64/ 5-Optimal-64 | 1-Solved-64 / 5-Solved-64 | SWC | Diversity | |
| Maze 15 x 15 | Complete-Trace | 495 | 94.6 / 90.1 | 96.7 /93.0 | 0.964 | 7.60 |
| Dualformer (slow) | 278 | 99.6 / 99.2 | 99.9 / 99.9 | 0.999 | 12.54 | |
| Mix-5% | 506 | 97.1 / 94.8 | 97.6 / 96.6 | 0.98 | 10.25 | |
| Mix-10% | 542 | 99.0 / 97.4 | 99.4 / 98.7 | 0.99 | 11.51 | |
| Mix-20% | 525 | 99.1 / 97.7 | 99.7 / 99.1 | 1.00 | 12.10 | |
| Mix-40% | 556 | 98.4 / 96.9 | 99.5 / 98.8 | 0.99 | 11.77 | |
| Mix-50% | 527 | 94.2 / 92.1 | 96.7 / 95.0 | 0.96 | 10.75 | |
| Mix-60% | 577 | 88.1 / 85.6 | 94.2 / 92.1 | 0.93 | 6.48 | |
| Mix-80% | 540 | 81.7 / 78.8 | 89.9 / 88.1 | 0.88 | 2.37 | |
| Maze 20 x 20 | Complete-Trace | 851 | 98.3 / 95.5 | 98.8 / 93.0 | 0.987 | 14.53 |
| Dualformer (slow) | 439 | 98.9 / 97.8 | 99.9 / 99.7 | 0.998 | 18.86 | |
| Mix-5% | 789 | 98.6 / 98.1 | 99.5 / 99.1 | 0.99 | 17.71 | |
| Mix-10% | 866 | 98.9 / 98.2 | 99.8 / 99.3 | 1.00 | 18.52 | |
| Mix-20% | 838 | 97.9 / 95.9 | 99.4 / 98.6 | 0.99 | 17.57 | |
| Mix-40% | 900 | 93.5 / 90.4 | 97.4 / 95.1 | 0.97 | 16.51 | |
| Mix-50% | 905 | 92.3 / 89.1 | 96.3 / 94.3 | 0.96 | 16.30 | |
| Mix-60% | 891 | 72.6 / 67.9 | 86.5 / 82.7 | 0.84 | 3.92 | |
| Mix-80% | 943 | 62.4 / 56.4 | 80.9 / 75.4 | 0.77 | 2.62 | |
| Maze 25 x 25 | Complete-Trace | 1208 | 95.2 / 85.7 | 97.0 / 90.4 | 0.968 | 18.85 |
| Dualformer (slow) | 589 | 99.9 / 97.2 | 99.7 / 99.3 | 0.997 | 25.05 | |
| Mix-5% | 1109 | 97.9 / 95.3 | 99.0 / 97.2 | 0.99 | 21.00 | |
| Mix-10% | 1264 | 98.7 / 96.3 | 99.5 / 98.0 | 0.99 | 22.33 | |
| Mix-20% | 1283 | 97.9 / 95.0 | 99.1 / 97.9 | 0.99 | 23.68 | |
| Mix-40% | 1278 | 90.8 / 87.4 | 96.3 / 93.9 | 0.96 | 22.30 | |
| Mix-50% | 1334 | 81.0 / 75.5 | 93.3 / 88.9 | 0.91 | 18.02 | |
| Mix-60% | 1266 | 61.0 / 53.4 | 80.3 / 74.0 | 0.77 | 4.63 | |
| Mix-80% | 1501 | 57.2 / 50.3 | 78.5 / 70.3 | 0.75 | 3.76 | |
| Maze 30 x 30 | Complete-Trace | 1538 | 93.3 / 82.4 | 95.9 / 88.1 | 0.964 | 7.60 |
| Dualformer (slow) | 854 | 97.6 / 93.2 | 99.5 / 98.2 | 0.993 | 25.77 | |
| Mix-5% | 2022 | 96.3 / 92.0 | 98.4 / 96.0 | 0.98 | 24.50 | |
| Mix-10% | 1720 | 97.7 / 95.0 | 99.2 / 98.2 | 0.99 | 27.17 | |
| Mix-20% | 1851 | 95.0 / 91.4 | 98.3 / 96.2 | 0.98 | 25.02 | |
| Mix-40% | 1854 | 64.3 / 56.2 | 83.1 / 76.0 | 0.81 | 12.60 | |
| Mix-50% | 1652 | 58.0 / 50.8 | 78.6 / 71.9 | 0.76 | 9.66 | |
| Mix-60% | 1983 | 43.0 / 35.4 | 68.4 / 58.1 | 0.65 | 3.21 | |
| Mix-80% | 1648 | 47.8 / 38.4 | 71.9 / 61.9 | 0.68 | 3.98 | |
| Sokoban | Complete-Trace | 3600 | 92.9 / 84.4 | 94.7 / 89.0 | 0.944 | 2.91 |
| Dualformer (slow) | 1482 | 94.5 / 87.6 | 97.4 / 94.1 | 0.97 | 4.66 | |
| Mix-5% | 3278 | 85.3 / 72.7 | 91.0 / 80.9 | 0.90 | 3.18 | |
| Mix-10% | 3402 | 88.6 / 77.2 | 94.1 / 87.9 | 0.93 | 4.07 | |
| Mix-20% | 3331 | 89.0 / 81.3 | 95.7 / 89.1 | 0.95 | 4.22 | |
| Mix-40% | 3294 | 92.7 / 86.1 | 97.1 / 93.1 | 0.96 | 4.14 | |
| Mix-50% | 3202 | 92.3 / 87.3 | 96.2 / 93.4 | 0.96 | 4.66 | |
| Mix-60% | 2594 | 91.1 / 83.2 | 96.4 / 91.0 | 0.96 | 4.48 |
| Maze 15x15 | Method | 1-Optimal-64 / 3-Optimal-64 | 1-Solved-64 / 3-Solved-64 | SWC | Diversity |
| Dualformer (fast) | 91.8 / 87.6 | 97.1 / 94.8 | 0.960 | 9.05 | |
| Mix-5% | 89.8 / 83.2 | 92.7 / 89.0 | 0.92 | 11.72 | |
| Mix-10% | 80.3 / 75.6 | 90.5 / 86.9 | 0.88 | 7.01 | |
| Mix-20% | 78.8 / 77.6 | 87.8 / 86.4 | 0.86 | 1.97 | |
| Mix-40% | 81.2 / 78.8 | 88.6 / 86.9 | 0.87 | 1.73 | |
| Mix-50% | 79.0 / 76.4 | 87.7 / 85.8 | 0.86 | 1.92 | |
| Mix-60% | 81.3 / 78.0 | 89.3 / 87.3 | 0.88 | 1.95 | |
| Mix-80% | 79.8 / 76.6 | 88.6 / 86.0 | 0.87 | 2.12 | |
| Solution-Only | 72.0 / 68.9 | 82.7 / 80.1 | 0.80 | 1.52 | |
| Maze 20x20 | Dualformer (fast) | 90.9 / 84.0 | 97.0 / 94.0 | 0.960 | 17.27 |
| Mix-5% | 86.2 / 74.8 | 94.0 / 87.7 | 0.93 | 15.92 | |
| Mix-10% | 88.6 / 81.4 | 95.2 / 91.7 | 0.94 | 16.04 | |
| Mix-20% | 65.2 / 60.5 | 81.7 / 77.1 | 0.79 | 2.72 | |
| Mix-40% | 63.2 / 59.0 | 80.4 / 77.2 | 0.78 | 2.65 | |
| Mix-50% | 63.3 / 57.9 | 79.6 / 76.0 | 0.77 | 2.39 | |
| Mix-60% | 66.8 / 63.6 | 82.4 / 80.1 | 0.80 | 2.50 | |
| Mix-80% | 56.5 / 51.8 | 74.5 / 70.2 | 0.71 | 2.13 | |
| Solution-Only | 56.3 / 52.0 | 71.9 / 67.5 | 0.69 | 1.52 | |
| Maze 25x25 | Dualformer (fast) | 83.9 / 72.9 | 95.5 / 90.6 | 0.940 | 21.23 |
| Mix-5% | 73.4 / 56.9 | 87.0 / 77.3 | 0.85 | 13.37 | |
| Mix-10% | 80.2 / 68.0 | 90.4 / 83.1 | 0.89 | 16.19 | |
| Mix-20% | 64.8 / 57.5 | 85.0 / 78.0 | 0.81 | 10.51 | |
| Mix-40% | 50.7 / 46.3 | 73.8 / 68.9 | 0.69 | 3.60 | |
| Mix-50% | 56.5 / 50.8 | 77.9 / 71.9 | 0.74 | 3.15 | |
| Mix-60% | 53.6 / 48.9 | 75.9 / 69.8 | 0.72 | 3.07 | |
| Mix-80% | 48.2 / 44.2 | 69.6 / 64.4 | 0.66 | 2.88 | |
| Solution-Only | 39.7 / 34.7 | 60.3 /55.4 | 0.57 | 1.9 | |
| Maze 30x30 | Dualformer (fast) | 80.0 / 66.0 | 91.8 / 85.7 | 0.906 | 18.23 |
| Mix-5% | 66.3 / 46.1 | 83.6 / 72.0 | 0.82 | 11.85 | |
| Mix-10% | 76.9 / 65.8 | 90.8 / 84.7 | 0.89 | 19.35 | |
| Mix-20% | 55.4 / 43.5 | 78.3 / 67.4 | 0.75 | 9.97 | |
| Mix-40% | 43.4 / 37.3 | 69.7 / 63.2 | 0.66 | 3.79 | |
| Mix-50% | 43.0 / 38.7 | 68.1 / 62.9 | 0.64 | 3.23 | |
| Mix-60% | 39.2 / 33.8 | 65.2 / 59.0 | 0.61 | 3.09 | |
| Mix-80% | 38.0 / 32.3 | 62.2 / 55.5 | 0.58 | 2.75 | |
| Solution-Only | 30.0 / 26.0 | 54.1 / 47.8 | 0.50 | 1.86 | |
| Sokoban | Dualformer (fast) | 97.3 / 94.4 | 94.8 / 90.0 | 0.970 | 4.92 |
| Mix-5% | 84.0 / 73.0 | 91.0 / 84.0 | 0.90 | 4.30 | |
| Mix-10% | 94.0 / 87.0 | 97.0 / 94.0 | 0.97 | 5.58 | |
| Mix-20% | 94.0 / 89.0 | 97.0 / 95.0 | 0.97 | 4.10 | |
| Mix-40% | 96.0 / 93.0 | 98.0 / 97.0 | 0.98 | 4.15 | |
| Mix-50% | 97.0 / 96.0 | 99.0 / 98.0 | 0.99 | 3.38 | |
| Mix-60% | 98.0 / 97.0 | 99.0 / 99.0 | 0.99 | 3.25 | |
| solution-only | 86.8 / 83.4 | 92.8 / 90.0 | 0.92 | 1.24 |
Appendix G Comparison with Different Randomization Strategies
In Table G.1 below, we show the probabilities of different dropping strategies we use for comparing different randomization strategies.
| Model | Probabilities of Dropping Strategies | |
| Maze | Dropping Level 1 | |
| Dropping Level 1 + 2 | ||
| Dropping Level 1 + 2 + 3 | ||
| Dualformer | ||
| Sokoban | Dropping Level 1 | |
| Dropping Level 1 + 2 | ||
| Dropping Level 1 + 2 + 3 | ||
| Dualformer |
Appendix H Details of the Math Reasoning Experiments
H.1 Solution Rewriting
H.1.1 Solution Rewriting Prompt
We use the following prompt template for rewriting the solutions via LLama-3.1-70B-Instruct.
H.1.2 Solution Rewriting Examples
Below, we provide an example of a math question along with four generated solutions that have been rewritten by the Llama-3.1-70B-Instruct model. It is evident from this example that the original math dataset lacks sufficiently detailed solution steps required to solve the question, such as the formula for a geometric series.
H.2 fine-tuning and Evaluation Prompts
We use the following prompt for fine-tuning both the Mistral and the Llama model, following Yu et al. (2023).
We use the following prompt for the slow mode evalaution.
For fast mode evaluation, we force the generation to directly output the final answer by adding the phrase “The answer is: ”.
H.3 Example Outputs
We provide a few example math questions and answers randomly sampled from our trained model, alongside the answers output by the baseline model. Within each answer box, on the left is the baseline model output (long and lengthy) and on the right is our method output (effective and efficient, colored in blue).
Appendix I LLM Generated Solutions for Maze
Readers might wonder whether modern state-of-the-art LLMs can effectively solve our maze problem. To test this, we randomly selected a maze problem and ask the o1-preview model to find the shortest path. O1-preview is the latest reasoning model by OpenAI which operates in slow mode: it spends more time thinking before they respond. It turns out that these problems are very challenging for LLMs. As illustrated in Figure I.1, the path suggested by o1-preview incorrectly traverses through the maze walls. In contrast, Dualformer correctly identifies one optimal path that follows through the maze without any errors.
Below, we also provide the exact prompt we used for o1 models, and their responses.